Line  Displaying Suppressing
Comment  continuation  output  output
Axiom  -- comment  input <cr>more input  input  input;
Derive  "comment"  input ∼<cr>more input  input  var:= input
MACSYMA /* comment */ input<cr>more input  input;  input$
MAPLE  # comment  input<cr>more input  input;  input:
Mathematica (* comment *) input<cr>more input  input  input;
MATLAB  % comment  input ..<cr>more input input  input;
MuPAD  # comment #  input<cr>more input  input;  input:
REDUCE  % comment  input<cr>more input  input;  input$
Loading a file  Timing a command  Quitting

Axiom  )read "file" )quiet  )set messages time on    )quit

Derive

MACSYMA load("file")$        showtime: all$           quit();
MAPLE  read(`file`):        readlib(showtime): on;   quit

Mathematica << file  Timing[command]         Quit[]
MATLAB  file  t0 = clock; command,    quit

                                   dt = etime(clock, t0)

MuPAD  read("file"):        time(command);          quit
REDUCE  in "file"$           on time;                 quit;

Previous

List  Defining a matrix  i π expression

Axiom  [1, 2]  m:= matrix([[1, 2],[3, 4]])  %i  %pi     %
Derive  [1, 2]  m:= [[1, 2],[3, 4]]          #i  pi
MACSYMA [1, 2]  m: matrix([1, 2],[3, 4]);    %i  %pi     %
MAPLE  [1, 2]  m:= array([[1, 2],[3, 4]]);  I   Pi      "

Mathematica {1, 2} m = {{1, 2},{3, 4}}  I   Pi      %

MATLAB  [1 2]   m = [1 2; 3 4]               i   pi      ans
MuPAD  [1, 2]  M:= Matrix(Integer):         I   PI      %

                      m:= M([[1, 2], [3, 4]]);

REDUCE  {1, 2} m:= mat((1, 2),(3, 4));      i   pi      ws

Table 1 Synonyms